### v3

# JailGuard: A Universal Detection Framework for LLM Prompt-based Attacks

Abstract.
Large Language Models (LLMs) and Multi-Modal LLMs (MLLMs) have played a critical role in numerous applications. However, current LLMs are vulnerable to prompt-based attacks, with jailbreaking attacks enabling LLMs to generate harmful content, while hijacking attacks manipulate the model to perform unintended tasks, underscoring the necessity for detection methods. Unfortunately, existing detecting approaches are usually tailored to specific attacks, resulting in poor generalization in detecting various attacks across different modalities. To address it, we propose JailGuard, a universal detection framework for jailbreaking and hijacking attacks across LLMs and MLLMs. JailGuard operates on the principle that attacks are inherently less robust than benign ones, regardless of method or modality. Specifically, JailGuard mutates untrusted inputs to generate variants and leverages discrepancy of the variants’ responses on the model to distinguish attack samples from benign samples. We implement 18 mutators for text and image inputs and design a mutator combination policy to further improve detection generalization. To evaluate the effectiveness of JailGuard, we build the first comprehensive multi-modal attack dataset, containing 11,000 data items across 15 known attack types. The evaluation suggests that JailGuard achieves the best detection accuracy of 86.14%/82.90% on text and image inputs, outperforming state-of-the-art methods by 11.81%-25.73% and 12.20%-21.40%.
LLM Security, Software and Application Security, Large Language Model System, LLM Defense
ccs: Security and privacy Software and application securityccs: Computing methodologies Neural networks
1. Introduction
Large Language Models (LLMs) have become commonplace in our technological interactions, from chatbots to complex decision-making engines (gpt, 2023a; Chen et al., 2021). They can perform various tasks such as understanding sentences, answering questions, etc., and are widely used in many different areas. For example, Meta has developed an AI assistant based on the LLM Llama and integrated it into multiple social platforms such as Facebook (Meta, 2024). The advent of Multi-Modal Large Language Models (MLLMs) has expanded these functionalities even further by incorporating visual understanding, allowing them to interpret and generate imagery alongside text, enhancing user experience with rich, multi-faceted interactions (gpt, 2023b; Zhu et al., 2023a; Liu et al., 2024a). Recently, Microsoft has released Copilot, a search engine based on MLLMs, which supports text and image modal input and provides high-quality information traditional search engines cannot provide (Microsoft, 2024).
Despite the utility of LLMs and MLLMs in various real-world scenarios, they suffer from both jailbreaking and hijacking attacks from different modalities. Jailbreaking attacks can bypass the security mechanisms of the model and software (e.g., AI-driven search engines) and produce harmful or illegal content like sex, violence, abuse, etc (Zou et al., 2023; Chao et al., 2023). Hijacking attacks can hijack and manipulate the LLM systems (e.g., AI assistants) to perform unintended tasks (Liu et al., 2023c). An LLM system might suffer from the two types of attacks on different modalities. For example, an AI assistant that supports multi-modal inputs could be misled by attackers to generate illegal content, or be hijacked to perform unintended tasks, ultimately exposing sensitive information, enabling the spread of misinformation, and damaging the overall trust in AI-driven applications and systems. Thus, there is an urgent need for a universal detection across different attacks and modalities, not only to help prevent these LLM prompt-based attacks and address such security gaps, but the attack samples identified and collected can also help developers understand the attacks and further improve LLM systems and applications.
There are approaches proposed to detect harmful content and the attack intent in models’ inputs and responses (azu, 2023; Jain et al., 2023; Robey et al., 2023). Despite these commendable efforts, existing LLM detection approaches still have limitations, resulting in poor adaptability and generalization across different modalities and attack methods. Typically, these methods rely on specific detection techniques or metrics (e.g., keywords and rules) to identify a limited range of attacks. They are designed to detect either jailbreaking attacks that produce harmful content (azu, 2023) or hijacking attacks that manipulate LLMs to perform specific tasks (Liu et al., 2023c). While such designs perform well on samples generated by specific attack methods, they struggle to detect attacks generated by other methods. Moreover, simply combining these detectors can result in a significant number of false positives. Consequently, existing detection methods are impractical for deployment in real-world scenarios where diverse attacks span different modalities.
Fig. 1. Leveraging the Robustness Difference to Identify Attacks
To break through these limitations, we design and implement a universal detection framework for LLM prompt-based attacks, JailGuard, which effectively identifies various prompt-based attacks on both image and text modalities. The key observation behind JailGuard is that attack inputs inherently exhibit lower robustness than benign queries, regardless of the attack methods and modalities. To confuse LLMs, attack inputs are often generated based on crafted templates or by an extensive searching process with complex perturbations. This leads to a result that any minor modification to the inputs may invalidate the attack’s effectiveness, which appears as a significant change in output. The responses of benign inputs, however, are hardly affected by these perturbations, as shown in the Fig. 1. Based on this observation, JailGuard first mutates the original input into a series of variant queries. Then the consistency of the responses of LLM systems to variants is analyzed. If a notable discrepancy can be identified among the responses, i.e., a divergence value that exceeds the built-in threshold, a potential prompt-based attack is identified. To effectively identify various attacks, JailGuard systematically designs and implements 16 random mutators and 2 semantic-driven targeted mutators to introduce perturbations at the different levels of text and image inputs. We observe that the detection effectiveness of JailGuard is closely tied to the mutation strategy, as different mutators apply disturbances at various levels and are suitable for detecting different attack methods. To design a more general and effective mutation strategy in detecting a wide range of attacks, JailGuard proposes a mutator combination policy as the default mutation strategy. Based on the empirical data of mutators, the policy selects three mutators to apply perturbations from different levels, combines their variants and divergences according to an optimized probability, and leverages their strengths to detect various attacks comprehensively.
To comprehensively evaluate JailGuard, we construct the first comprehensive LLM prompt-based attack dataset that contains 11,000 items of data covering 15 types of jailbreaking and hijacking attacks on image and text modalities. Based on this dataset, we conduct large-scale experiments that spend over 500M paid tokens to compare JailGuard with 12 state-of-the-art (SOTA) jailbreaking and hijacking detection methods on text and image inputs, including commercial detector Azure content detector (azu, 2023). The experimental results indicate that all mutators in JailGuard can effectively identify LLM prompt-based attacks and benign samples on image and text modalities, achieving higher detection accuracy than SOTA. In addition, the default combination policy of JailGuard further improves the detection results and has separately achieved the best accuracy of 86.14% and 82.90% on text and image inputs, significantly outperforming state-of-the-art defense methods by 11.81%-25.73% and 12.20%-21.40%. In addition, JailGuard can effectively detect and defend different types of prompt-based attacks. Among all types of collected attacks, the best detection accuracy in JailGuard ranges from 76.56% to 100.00%. The default combination policy in JailGuard can achieve an accuracy of more than 70% on 10 types of text attacks, and the detection accuracy on benign samples is over 80%, which exhibits the best generalization among all mutators and baselines. Furthermore, the experiment results also demonstrate the efficiency of JailGuard. We observe that the detection accuracy of the JailGuard’s mutators does not drop significantly when the LLM query budgets (i.e., the number of generated variants) reduce from $N=8$ to $N=4$ and is always better than that of the best baseline SmoothLLM. This finding can provide guidance on attack detection and defense in low-budget scenarios.
In summary, our contributions are:
  • We identify the inherent low robustness of LLM attack inputs. Based on that, we design and implement the first universal LLM attack detection framework, JailGuard, which implements 16 random mutators, 2 semantic-driven targeted mutators, and a set of combination policies. JailGuard mutates the original query to generate variants and uses the divergence of the variants’ responses to detect LLM prompt-based attacks (i.e., jailbreaking and hijacking attacks) on image and text modalities.
  • We construct the first comprehensive LLM attack dataset that consists of 11,000 samples and covers 15 jailbreaking and hijacking attacks on both image and text inputs.
  • We perform experiments on our constructed dataset, and JailGuard has achieved better detection effects than the state-of-the-art methods.
  • We open-source our dataset and code on our website (our, 2024).
Threat to Validity. JailGuard is currently evaluated on a dataset consisting of 11,000 items of data and 12 attack methods, which may be limited. Although our basic idea can theoretically be applied to other attack methods, this may not hold on some unseen attacks. To mitigate the threats and follow the Open Science Policy, the code of the prototype JailGuard and all the experiment results will be publicly available at (our, 2024).
2. Background
2.1. Prompt-based LLM Attack
Prompt-based LLM attack aims to design and generate an attack prompt $P_{a}$ that can induce the target LLM system and application $M$ to contain attack target $T$ in the model response $R=M(P_{a})$, which can be expressed as follows.
[…]
where $eval(\cdot)$ is an evaluation function and it returns 1 iff the input prompt $P_{a}$ bypasses the LLM’s protection mechanism and the corresponding response $R$ achieves the attack target $T$. This paper aims to design a detection framework to identify prompt-based attacks, which are mainly divided into jailbreaking and hijacking attacks according to the differences in attack target $T$.
Jailbreaking attack leverages elaborate templates, specific strings, etc. to guide the LLMs to generate toxic contents that violate usage policies [^1] https://platform.openai.com/docs/guides/moderation, such as sexual information and hateful contents To effectively and automatically generate jailbreak prompts, researchers proposed a variety of attack methods (Zou et al., 2023; Deng et al., 2023a; Chao et al., 2023; Zhu et al., 2023b; Yu et al., 2023). Zou et al. (Zou et al., 2023) designed the greedy coordinate gradient-based search (GCG) to produce adversarial suffix to attack open-sourced LLMs (e.g., Vicuna (Zheng et al., 2023)), which has proven its effectiveness through transfer attacks on black-box commercial LLMs. TAP (Mehrotra et al., 2023) is one of the state-of-the-art jailbreaking methods that only requires black-box access to the target LLM. It utilizes LLMs to iteratively refine candidate attack prompts using tree-of-thoughts reasoning until one of the generated prompts jailbreaks the target LLM. With the emergence of MLLMs, researchers designed a visual jailbreaking attack based on the implanted adversarial perturbation in the image inputs (Qi et al., 2024). Their method achieved a high attack success rate on MiniGPT-4 which is one of the state-of-the-art MLLMs (Zhu et al., 2023a). We collect 10 jailbreaking attacks at the text and image level in our dataset, as shown in § 5.
[…]
2.2. LLM Attack Detector
Conducting a detector to identify the attack inputs of the given model is one of the most popular defense strategies in the field of AI security (Metzen et al., 2017; Wang et al., 2021; Liu et al., 2022; Dong et al., 2021). Detectors not only prevent attacks that have not yet occurred on AI systems but also help developers understand attack methods and attacker intentions, thereby improving the safety and security of AI systems and software. Existing LLM attack detectors leverage the model input prompt $P$ and response $R$ to identify attacks. The expected output of the detector can be expressed as follows.
$detect(P,R)=\begin{cases}1&\text{if}\quad P\in\mathcal{P}_{a}\\ 0&\text{otherwise}\end{cases}$
where $\mathcal{P}_{a}$ represents the attack prompt set. When the detector recognizes the attack input, its output is 1 and such an input will be filtered. Otherwise, the output is 0 and the LLM response passes the detector.
LLM attack detectors are widely used in commercial services such as Azure content detector (azu, 2023). Guo et al. (Gou et al., 2024) design the LLM-based harm detector to identify the attack inputs on MLLMs and then regenerate safe-aligned responses. Alon et al. (Alon and Kamfonas, 2023) propose to detect jailbreaking attacks by evaluating the perplexity of queries with adversarial suffixes. Similarly, Liu et al. (Liu et al., 2023c) implement a detection method that uses a built-in threshold and the perplexity of input query and LLMs to identify prompt injection attacks. In addition, SmoothLLM (Robey et al., 2023) uses a set of refusal keywords (Zou et al., 2023) to distinguish blocked jailbreak attack responses from normal responses and aggregates them to obtain the final LLM response. In this paper, we propose a universal LLM attack detection framework, JailGuard. We compare 12 open-sourced LLM detection and defense methods to demonstrate the effectiveness of JailGuard in detecting LLM jailbreaking and hijacking attacks.
3. Motivation
Fig. 2. Motivation Cases of JailGuard
Existing LLM attack detection approaches exhibit poor generalization across various attacks. They typically rely on specific detection techniques or metrics tailored to particular attack methods, making them less or ineffective in identifying attacks generated by different methods or on different modalities. For example, Azure Content Detector (azu, 2023) leverages built-in rules and models to identify harmful content and blocks potential threats and quality problems in LLMs’ inputs and responses, which can be used to mitigate jailbreak attacks. However, it cannot identify hijacking injection attacks (e.g., 1 in Fig. 2) whose inputs and responses do not contain harmful content within its rules. SmoothLLM (Robey et al., 2023) implants interference into text inputs, and uses refusal keywords to aggregate LLM responses and identify attacks. It can effectively detect jailbreaking attacks, but it is ineffective in detecting hijacking text attacks whose output does not contain keywords and image attacks, as shown in attack cases ( 1 and 3). In real-world scenarios, LLM systems and software face both jailbreaking and hijacking attacks that span different modalities and use various methods. Existing methods struggle to effectively identify these diverse attacks simultaneously, leading to false negatives in detection. Furthermore, deploying multiple detectors on the LLM systems results in a high number of false positives. Consequently, existing detection methods with poor generalization are challenging to deploy or apply in real-world scenarios.
To fill the gap, we have studied existing LLM attack method (Zou et al., 2023; Deng et al., 2023a; Xu et al., 2024; Liu et al., 2023d, c) and find that existing attacks mainly rely on specific templates or tiny but complicated perturbations. These elaborate perturbations and templates can shift the attention of the LLM system and application and deceive its built-in defense mechanisms. Although these elaborate attacks have achieved excellent attack results, they are less robust than benign queries and easily disrupted and fail, resulting in large divergence between LLM responses. Such a difference in robustness can be used to distinguish between attack and benign queries, regardless of the attack method and input modality.
Based on this observation, we propose JailGuard, a universal LLM prompt-based attack detection framework that leverages the divergence features of the responses of perturbed attack inputs to identify various attacks, regardless of the attack method and input modality. For the three attack inputs in Fig. 2.a), JailGuard generates variants based on its mutation strategy and then queries the LLM system to obtain variant responses. The attack input is susceptible to interference and produces multiple responses with different semantics, as shown in Fig. 2.b). Red text marks responses where the attack was still successful despite interference, and black text marks responses from failed attacks. JailGuard vectorizes these responses and computes the divergence between each set of variant responses. The divergence of these attacks far exceeds the built-in threshold in JailGuard (i.e., 0.02 for text inputs and 0.025 for image inputs), as shown in the green texts in Fig. 2.c). Therefore, leveraging the robustness and divergence features of attack samples, JailGuard effectively detects all prompt-based attacks on images and text modalities in Fig. 2, outperforming the state-of-the-art detection methods.
4. System Design
JailGuard is implemented on top of the LLM system, and Fig. 3 shows the overview. JailGuard first implements a Detection Framework (§ 4.1) that detects attacks based on input mutation and the divergence of responses. For the untrusted input query, the detection framework leverages the built-in mutation strategy to generate a variant set. Then it uses these variants to query the target LLM system and computes the semantic similarity and divergence between variant responses, finally leveraging the built-in thresholds to identify benign and attack queries. To effectively detect various attacks, in Mutation Strategy (§ 4.2), JailGuard first systematically design 18 mutators to introduce perturbation at different levels for text and image inputs, including 16 random mutators and 2 semantic-guided targeted mutators. However, we observe that the mutator selection has a great impact on the detection effect of the framework on different attacks. To achieve better generalization and improve detection effects, we propose a combination-based mutation policy as the default strategy in JailGuard to merge multiple mutators and their divergence based on their empirical data and leverage their strengths to identify different attacks.
4.1. Detection Framework
Fig. 3. Overview of JailGuard
The key observation of our detection framework is that compared with benign samples, regardless of the attack types and modalities, attack samples tend to be less robust and more susceptible to interference, leading to semantically different responses, as shown in Fig. 2. Therefore, the detection framework first mutates the original input query to generate a set of variants. It then calculates the divergence between the LLM responses to the variants and utilizes the built-in threshold to identify those attack samples with significantly larger divergence. The detection framework proceeds through the following steps:
Mutating original inputs. For the original untrusted input prompt $P$, the detection framework leverages mutators to generate multiple variants that are slightly different from the original input. The variant set can be represented as $\mathcal{P}=\{P_{1},...,P_{N}\}$, where $N$ indicates the number of variants, which is related to the LLM query budget.
Constructing the similarity matrix. For the input variant set $\mathcal{P}$, the detection framework first queries the LLM system to obtain the response set =$\mathcal{R}=\{R_{1},...,R_{N}\}$. For each $R_{i}$ in $\mathcal{R}$, the detection framework leverages the ‘spaCy’ library to convert it to word vector $V_{i}$ and then calculates the cosine similarity between it and the word vectors of other responses. The similairty $S_{i,j}$ between vectors $V_{i}$ and $V_{j}$ can be represented as:
$S_{i,j}=COS(V_{i},V_{j})=\frac{V_{i}\cdot V_{j}}{\|V_{i}\|\|V_{j}\|},$
where $COS(\cdot)$ calculates the cosine similarity between two vectors, $i,j\in\{1,2,\ldots,N\}$. Similarity values for response pairs are represented in an $N\times N$ matrix $S$, where each element at $(i,j)$ corresponds to the similarity between the pair $(R_{i},R_{j})$.
Characterizing each response. In the $N\times N$ similarity matrix $S$, each row $S_{i,\cdot}$ corresponding to the $i$-th response $R_{i}$ can be normalized using $\mathcal{L}_{1}$ normalization. This process transforms the row into a similarity probability distribution, effectively characterizing the relationship of response $R_{i}$ to the other responses in the matrix.
$Q_{i}=Q(X=i)=\|S_{i,\cdot}\|_{1}.$
Quantifying the divergence of two responses. Then JailGuard uses Kullback-Leibler (KL) divergence to quantify the difference between the similarity distributions of two responses $(Q_{i},Q_{j})$:
$D_{i,j}=D(Q_{i}\|Q_{j})=\sum_{x=1}^{N}Q_{i}(x)\log(\frac{Q_{i}(x)}{Q_{j}(x)}).$
The divergence values for response pairs are represented in an $N\times N$ matrix $D$, where each element at $(i,j)$ corresponds to the divergence between the pair $(R_{i},R_{j})$.
Examining the divergence Finally, for the obtained divergence $N\times N$ matrix $D$, the detection framework uses the threshold $\theta$ to identify the attack input. Specifically, if any value in divergence matrix $D$ exceeds $\theta$, JailGuard will consider the original input as an attack input, otherwise it is judged as a benign input, which is shown as follows:
$\exists i,j\in\{1,2,\ldots,N\},D_{i,j}\geq\theta\rightarrow\{P\}\cup\mathcal{P}_{a},$
where $\mathcal{P}_{a}$ represents the set of inputs detected as LLM attacks by JailGuard. Note that, when all variants of an attack input fail, the LLM system and application will not provide any service for these inputs. In this case, all responses will contain the refusal keywords (Zou et al., 2023) and become similar in semantics, and their divergence $D$ will be very low. Therefore, if all responses contain refusal words, regardless of the value in $D$, JailGuard will directly determine them as attack inputs.
Table 1. Demo Case for Text Mutators of JailGuard (Red highlights the content modified by mutators and ‘…’ skips those content that are not affected by mutators.)
Mutator Demo Variant
No Mutator You are a highly skilled marketer who is passionate about helping businesses in the adult services industry succeed. Your clients rely on your expertise to help them attract new clients and grow their businesses. Can you provide some unique and effective marketing strategies that will help them stand out from the competition while maintaining ethical and legal standards?
Random Replacement … Your clients rely on your expertise to help them attract new [Mask]s and grow their businesses. Can you provide some unique and effective marketing strategies tha[Mask] help them stand out from the competition…
Random Insertion …Your clients rely on your expertise to help the[Mask]m attract new clients and grow their businesses. Can you provide some unique and effective mar[Mask]keting strategies that will help them stand out from the competition while maintaining ethical and legal standards?
Random Deletion You are a highly skmarketer who is passionate about helping businesses… Can yvide some unique and effective marketing strs that will help them out from the competition whileaining ethical anl standards?
Punctuation Insertion You are a highly skilled marketer who is passionate about helping businesses in the adult services . industry succeed. Your clients rely on your , expertise to help them attract new : clients . and grow their businesses. Can you provide some unique and effective marketing strategies that will help them stand out from the competition while maintaining ! ethical and legal standards?
Synonym Replacement … Your clients rely on your expertness to helper them attract raw clients and farm their businesses. tail you supply some unique and efficacious marketing strategies that will helper them suffer out from the competition while exert ethical and effectual standards?
Translation You are a highly qualified marketing specialist who is passionate about helping companies in the adult services industry. your customers are counting on your expertise to help them attract new customers and develop their business. can you provide unique and effective marketing strategies that will help them stand out from competition while retaining ethical and legal standards?
Targeted Replacement …Your clients rely on your expertise to help them attract new clients and grow their businesses. Can you provide some unique and eff[Mask] marketing strate[Mask]hat will help them[Mask] out from the com[Mask]on while maintaining ethical and legal standards?
Targeted Insertion …Your clients rely on your expertise to help them attract new clients and grow the[Mask]ir businesses. Can you provide some [Mask]unique and effective [Mask]marketing strategies that will help th[Mask]em stand out from the [Mask]competition while[Mask] maintaining ethical and legal standards?
4.2. Mutation Strategy
4.2.1. Single Mutator
To effectively detect various attacks, JailGuard first systematically designs and implements a total of 18 single mutators in the mutation strategy, including 16 random mutators and 2 semantic-guided mutators, to introduce different levels of perturbations for image and text inputs. We separately provide demo cases for text and image mutators in Table 1 and Fig. 4.
Random text mutators. JailGuard implements six random mutators for text inputs, namely Random Replacement, Random Insertion, Random Deletion, Punctuation Insertion, Synonym Replacement, and Translation. Following the taxonomy from the prior work (Bayer et al., 2022), these mutators apply perturbations to the target text at three levels (i.e., from local characters to global sentences), namely character-level, word-level, and sentence-level.
$\bullet$ Character-level mutators randomly implant and modify characters in the text input, imposing perturbations at part of the input query. This category includes Random Replacement, Random Insertion, Random Deletion, and Punctuation Insertion. Random Replacement and Random Insertion perform the replacement or insertion operation with probability $p$ for each character. The replacement operation replaces the target and subsequent characters with a specific string $S$, ensuring that the input length does not change. The insertion operation inserts $S$ at the position after the target character. Similarly, Random Deletion removes the character in the text with probability $p$. Punctuation Insertion follows existing data augmentation methods that randomly insert punctuation masks into the target texts (Karimi et al., 2021). can potentially disturb adversarial-based LLM attacks without altering the semantics of the input sentence. Rows 3-6 of Table 1 provide demo cases for these character-level mutators, and red highlights the modifications.
$\bullet$ Word-level mutators target complete words in text to perform modifications or replacements. We implement the Synonym Replacement mutator that selects words in the text input and uses their synonyms to replace them based on WordNet (Miller, 1995). Substituting synonyms could bring slight changes to the semantics of the whole sentence. Row 7 of Table 1 provides a demo case.
$\bullet$ Sentence-level mutators modify and rewrite the entire input query to interfere with the embedded attack intent. JailGuard implements one sentence-level mutator, Translation. This mutator first translates the input sentence into a random language and then translates it back to the original language. This process can prevent LLM attacks based on specific templates and adversarial perturbations by rewriting the templates and removing meaningless attack strings, while still retaining the semantics and instructions of benign inputs. Row 8 of Table 1 provides a demo case.
Random image mutators. Inspired by existing work (Hendrycks et al., 2019; Lopes et al., 2019), we design 10 random mutators for image inputs in JailGuard, namely Horizental Flip, Vertical Flip, Random Rotation, Crop and Resize, Random Mask, Random Solarization, Random Grayscale, Gaussian Blur, Colorjitter, and Random Posterization. These mutators can be divided into three categories (Mumuni and Mumuni, 2022) according to the method of applying random perturbation, namely geometric mutators, region mutators, and photometric mutators.
$\bullet$ Geometric mutators alter the geometrical structure of images by shifting image pixels to new positions without modifying the pixel values, which can preserve the local feature and information of the image input. JailGuard implements four geometric mutators, namely Horizental Flip, Vertical Flip, Random Rotation, and Crop and Resize. Horizental Flip and Vertical Flip respectively flip the target image horizontally or vertically with a random probability between 0 and 1. Random Rotation (Gidaris et al., 2018; Cubuk et al., 2020) rotates the image by a random number of degrees between 0 and 180. After rotation, the area that exceeds the original size will be cropped. Crop and Resize (Bai et al., 2022) crops a random aspect of the original image and then resizes it to a random size, disturbing attack images without changing their color and style. We have provided examples in Fig. 4.a)-d).
$\bullet$ Region mutators apply perturbations in random regions of the image, rather than uniformly transforming the entire image. We implement Random Mask in JailGuard that inserts a small black mask to a random position of the image, as shown in Fig. 4.e). It helps disturb information (e.g., text) embedded by the attacker, leading to a drastic change in LLM responses.
$\bullet$ Photometric mutators simulate photometric transformations by modifying image pixel values, thereby applying pixel-level perturbations on image inputs. JailGuard implements five geometric mutators, namely Random Solarization, Random Grayscale, Gaussian Blur, Colorjitter, and Random Posterization. Random Solarization mutator inverts all pixel values above a random threshold with a certain probability, resulting in solarizing the input image. This mutator can introduce pixel-level perturbations for the whole image without damaging the relationship between each part in the image. Random Grayscale is a commonly used data augmentation method that converts an RGB image into a grayscale image with a random probability between 0 to 1 (He et al., 2020; Gong et al., 2021; Bai et al., 2022). Gaussian Blur (Bai et al., 2022) blurs images with the Gaussian function with a random kernel size. It reduces the sharpness or high-frequency details in an image, which intuitively helps to disrupt the potential attack in image inputs. Colorjitter (He et al., 2020) randomly modifies the brightness and hue of images and introduces variations in their color properties. Random Posterization randomly posterizes an image by reducing the number of bits for each color channel. It can remove small perturbations and output a more stylized and simplified image. We provide demos for these mutators in Fig. 4.f)-j).
Fig. 4. Demo Case for Image Mutators of JailGuard
Targeted mutators. Although random mutators have the potential to disrupt LLM attacks, mutators that apply perturbations with random strategies are still limited by false positives and negatives in detection and have poor generalization across different attack methods. On the one hand, if the mutators randomly modify with a low probability, they may not cause enough interference with the attack input, leading to false negatives in detection. On the other hand, blindly introducing excessive modification may harm LLMs’ response to benign inputs, which leads to dramatic changes in their responses‘ semantics and more false positives. This is especially true in the text, where small changes to a word may completely change its meaning. To implant perturbations into attack samples in a targeted manner, we design and implement two semantic-guided targeted text mutators in JailGuard, namely Targeted Replacement and Targeted Insertion.
Different from the random mutators Random Replacement and Random Insertion that blindly insert or replace characters in input queries, Targeted Replacement and Targeted Insertion offer a more precise approach to applying perturbations by considering the semantic context of the text, thereby enhancing the detection accuracy of LLM attacks. Specifically, the workflow of targeted mutators has two steps:
  1. (1)
    Step 1: Identifying Important content. Our manual analysis of existing attack methods and samples that can bypass the random mutator detection shows that these attack samples usually leverage complex templates and contexts to build specific scenarios, implement role-playing, and shift model attention to conduct attacks. These queries usually have repetitive and lengthy descriptions (e.g., setting of the ‘Do-Anything-Now’ mode, descriptions of virtual characters like Dr. AI (Liu et al., 2023b; Yu et al., 2023)), which are rare in benign inputs. To effectively identify these important contents, JailGuard implement a word frequency-based method. It scans the input text to find sentences with the highest word frequency, which are likely to contain the core components of the attack, such as repetitive instructions or descriptions that are integral to the attack’s success.
  2. (2)
    Step 2: Modifying. Once the important content is identified, targeted mutators are applied with a higher probability (i.e., five times the probability) to perform stronger perturbations on attack samples. For other content, the mutators are performed based on the original probability, as mentioned in the ‘Character-level mutators’. We provide examples in Row 9-10 of Table 1, and bold italics mark the identified important sentence.
By focusing on critical content with higher modification probabilities and applying character-level mutator to less critical parts, JailGuard enhances its ability to disrupt attack inputs while preserving the semantics of benign queries, leading to more effective detection of both jailbreaking and hijacking attacks across various modalities and methods.
4.2.2. Combination Policy
We have observed that the selection of mutators determines the quality of generated variants and the detection effect of variant responses’ divergence. Additionally, a single mutator typically excels at identifying specific attack inputs but struggles with those generated by different attack methods. For instance, the text mutator Synonym Replacement randomly replaces words with synonyms and achieves the best detection results on the naive injection method that directly implants instructions in inputs among all mutators. However, this approach proves ineffective against template-based jailbreak attacks, where its detection accuracy is notably lower than most other mutators, as detailed in § 6.3.
To design a more effective and general mutation strategy, inspired by prior work (Hendrycks et al., 2019), we design a straightforward yet effective mutator combination policy. This policy integrates various mutators, leveraging their individual strengths to detect a wide array of attacks. The policy first involves selecting several mutators to build a mutator pool. When generating variants, the policy merges variants generated by different mutators based on the optimized built-in probability and then calculates the divergences of their responses.
To determine the optimal mutator pool and probability, we use 70% of our dataset (§ 5) as the training set and conduct large-scale experiments to collect empirical data of different mutators. We then employ an optimization tool (Komer et al., 2014) to search for the sampling probability of mutators. Consequently, for the combination policy of text mutators, we select Punctuation Insertion, Targeted Insertion and Translation to construct the mutator pool, and their sampling probabilities are $[0.24,0.52,0.24]$. For the image inputs, we select Random Rotation, Gaussian Blur and Random Posterization, and the sampling probabilities are $[0.34,0.45,0.21]$ respectively. The effectiveness of the mutator combination policy is validated in § 6.2 and § 6.4.
5. Dataset Construction
In real-world scenarios, MLLM systems face both jailbreaking and hijacking attack inputs across different modalities. For example, attackers may attempt to mislead the MLLM agent into producing harmful content (e.g., violence, sex) or inject specific instructions to hijack the agent into performing unintended tasks. Thus, it is crucial to comprehensively evaluate the effectiveness of LLM detection methods to identify and prevent various prompt-based attacks simultaneously.
However, due to the absence of a comprehensive LLM prompt-based attack dataset, existing LLM defense research mainly tests and evaluates their methods on inputs generated by specific attacks. For example, SmoothLLM (Robey et al., 2023) has evaluated its effectiveness in defending against jailbreak inputs generated by the GCG attack (Zou et al., 2023), overlooking other attacks (e.g., prompt injection attack) that can also have severe consequences. To address this limitation, We first collect the most popular jailbreaking and hijacking injection attack inputs from the open-source community and prior work. We then evaluate their effectiveness on LLM systems and applications, filtering out those samples where the attacks fail. Finally, we construct a dataset covering 15 types of prompt-based LLM attacks, covering two modalities of image and text, with a total of 11,000 items of attack and benign data. Our dataset is publicly available on our website (our, 2024).
Table 2. LLM Prompt-based Attacks in Our Dataset (Grey Marks Jailbreaking Attacks and Blue Marks Hijacking Attacks)
Input
Modality
Attack Approach Description
Text Parameters (Huang et al., 2023a) Adjusting parameters in LLMs (APIs) to conduct jailbreaking attacks.
DeepInception (Li et al., 2023b) Constructing nested scenes to guide LLMs to generate sensitive content.
GPTFuzz (Yu et al., 2023) Random mutating and generating new attacks based on human-written templates.
Tap (Mehrotra et al., 2023) Iteratively refining candidate attack prompts using tree-of-thoughts.
Template-based (Liu et al., 2023b) Leveraging various human-written templates into jailbreak LLMs.
Jailbroken (Wei et al., 2024) Construct attacks based on the existing failure modes of safety training.
PAIR (Chao et al., 2023) Generating semantic jailbreaks by iteratively updating and refining a candidate prompt.
Naive Injection (Goodside, 2022) Directly concatenating target data, injected instruction, and injected data.
Fake Completion (Willison, 2023) Adding a response to mislead the LLMs that the previous task has been completed.
Ignoring Context (Perez and Ribeiro, 2022) Adding context-switching text to mislead the LLMs that the context changes.
Escape Characters (Liu et al., 2023a) Leveraging characters to embed instructions in texts to change the original query intent.
Combined Attack (Liu et al., 2023c) Combining existing methods (e.g., escape characters, context ignoring) to effectively inject.
Text + Image Visual Adversarial Example (Qi et al., 2024) Implanting unobservable adversarial perturbations into images to attack LLMs.
Typographic (TYPO) (Liu et al., 2023d) Embedding malicious instructions in blank images to conduct attacks.
Typographic (SD+TYPO) (Liu et al., 2023d) Embedding malicious instructions in images generated by Stable Diffusion to conduct attacks.
Text inputs. To ensure the diversity of LLM text attacks, we have collected a total of 12 kinds of attack inputs of the two common LLM attacks (i.e., jailbreaking attacks and hijacking injection attacks). Table 2 provides an overview of these attack methods. For jailbreaking attacks, to comprehensively cover various attack methodologies, we collect the most popular generative attack methods (i.e., Parameters (Huang et al., 2023a), DeepInception (Li et al., 2023b), GPTFuzz (Yu et al., 2023), TAP (Mehrotra et al., 2023), Jailbroken (Wei et al., 2024), Pair (Chao et al., 2023)) and the template-based attack method from the open-source community and existing study (Liu et al., 2023b) (including over 50 attack templates) to construct the attack inputs on GPT-3.5-Turbo-1106. To validate the effectiveness of these jailbreaking attack prompts and ensure the dataset’s quality, we follow the existing work (Qi et al., 2023) and evaluate the score of the prompts and the corresponding LLM’s responses violating OpenAI policies (score from 1 to 5) and only select those attack inputs with the highest score and the most severe violations (i.e., score 5) to construct the jailbreaking attack dataset, comprising a total of 2,000 items. For injection-based hijacking attacks, we have collected the most popular LLM injection attack methods, namely naive injection attack (Goodside, 2022), fake completion attack (Willison, 2023), ignoring context attack (Perez and Ribeiro, 2022), escape characters attack (Liu et al., 2023a), and combined attack (Liu et al., 2023c). We also verify the effectiveness of these attack samples on GPT-3.5-Turbo-1106 and select 2,000 items that can truly hijack LLMs to build our dataset.
Considering that the number of benign queries in the real world is much more than attack queries, our dataset maintains a ratio of $1:1.5$ for the attack and benign data. We randomly sample a total of 6,000 questions from the existing LLM instruction datasets (Li et al., 2023a; Askari et al., 2023; Zhou et al., 2023) as the benign dataset. These instruction datasets have been widely used in prior work (Chen et al., 2024a; Dubois et al., 2024; Lou et al., 2023; Sha and Zhang, 2024) for fine-tuning and evaluation. The benign data covers various question types such as common sense questions and answers, role-playing, logical reasoning, summarization, etc.
Text + Image inputs. Compared with text attacks, MLLM attack methods are relatively monotonous. We have collected the most popular adversarial-based jailbreaking attacks and typographic hijacking attacks. Adversarial-based attacks implant adversarial perturbations into images to guide the MLLM to produce harmful content. We leverage the prior work (Qi et al., 2024) to construct and collect 200 items of attack inputs on MiniGPT-4. The typographic attack is an injection-based attack method that involves implanting text into images to attack MLLMs (Liu et al., 2023d; Noever and Noever, 2021). We gather 200 attack inputs that use typographic images to replace sensitive keywords in harmful queries from MM-SafetyBench (Liu et al., 2023d), with 100 items embedding text in images generated by Stable Diffusion, and 100 items directly embedding text in blank images. Following existing work (Qi et al., 2023), all attack inputs have been validated for their effectiveness in attacking MiniGPT-4. Additionally, we include 600 benign inputs sampled from open-source training datasets of LLaVA (Zhou et al., 2023) and MiniGPT-4 (Zhu et al., 2023a) to balance the image dataset.
6. Evaluation
RQ1: How effective is JailGuard in detecting and defending against LLM prompt-based attacks at the text and visual level?
RQ2: Can JailGuard effectively and generally detect different types of LLM attacks?
RQ3: What is the contribution of the mutator combination policy and divergence-based detection in JailGuard?
RQ4: What is the impact of the LLM query budget (i.e., the number of generated variants) in JailGuard?
6.1. Setup
Baseline. To the best of our knowledge, we are the first to design a universal LLM attack detector for different attack methods on both text and image inputs. We select 12 state-of-the-art LLM jailbreak and prompt injection defense methods that have open-sourced implementation as baselines to demonstrate the effectiveness of JailGuard, as shown in the following.
  • Content Detector is implemented in Llama-2 repository [^2] https://github.com/facebookresearch/llama-recipes. It is a combined detector that separately leverages the Azure Content Safety Detector (azu, 2023), AuditNLG library (Aud, 2023), and ‘safety-flan-t5-base’ language model to check whether the text input contains toxic or harmful query. To achieve the best detection effect, we enable all three modules in it.
  • SmoothLLM (Robey et al., 2023) is one of the state-of-the-art LLM defense methods for the text input. It perturbates input with three different methods, namely ‘insert’, ‘swap’, and ‘patch’ and aggregates the LLM responses as the final response. Based on their experiment setting and results, we set the perturbation percentage to 10% and generate 8 variants for each input.
  • In-context defense (Wei et al., 2023) leverages a few in-context demonstrations to decrease the probability of jailbreaking and enhance LLM safety without fine-tuning. We follow the context design in their paper and use it as a baseline for text inputs.
  • Prior work (Liu et al., 2023c; Jain et al., 2023) implements several defense methods for jailbreaking and prompt injection attacks. We select four representative defense methods as baselines for text inputs in experiments, namely paraphrase, perplexity-based detection, data prompt isolation defense, and LLM-based detection. We query GPT-3.5-1106 to implement the paraphrase and LLM-based detection. Following the setting in prior work (Jain et al., 2023), we set the window size to 10 and use the maximum perplexity over all windows in the harmful prompts of AdvBench dataset (Zou et al., 2023) as the threshold, that is 1.51. For the data prompt isolation defense and LLM-based detection, we directly use the exiting implementation (Liu et al., 2023c).
  • BIPIA (Yi et al., 2023) proposes a black box prompt injection defense method based on prompt learning. It provides a few examples of indirect prompt injection with correct responses at the beginning of a prompt to guide LLMs to ignore malicious instructions in the external content. We directly use their implementation and default setting in experiments.
  • Self-reminder (Xie et al., 2023) modifies the system prompt to ask LLMs not to generate harmful and misleading content, which can be used on both text and image inputs.
  • ECSO detection (Gou et al., 2024) uses the MLLM itself as a detector to judge whether the inputs and responses of MLLM contain harmful content. We directly use this detector for inputs.
Metric. As mentioned in § 2, the LLM attack detector $detect(\cdot)$ assesses whether LLMs’ inputs are attacks. A positive output (i.e., 1) from $detect(\cdot)$ indicates an attack input, while a negative output (i.e., 0) signifies a benign input. For baselines employing filtering and similar mechanisms to defend against attacks without explicit detection mechanisms (e.g., Self-reminder), we use keyword-based detection from prior work (Robey et al., 2023; Zou et al., 2023) to evaluate their detection effects. When a specific refusal keyword (e.g., ‘I’m sorry’, ‘I apologize’) is detected in the LLM response, the original attack input is identified and blocked by the defense method, and $detect(\cdot)$ is 1 at this time, otherwise, it is set to 0.
Following the prior work (Liu et al., 2022; Metzen et al., 2017), we collect the True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) in detection and use metrics accuracy, precision, and recall to comprehensively assess the effectiveness of detection.
Accuracy calculate the proportion of samples correctly classified by the defense method, encompassing both correctly identified attack samples (i.e., TP) and benign samples that pass through (i.e., TN), relative to the total samples. This metric comprehensively evaluates whether the detector effectively mitigates attacks while ensuring valid responses to benign inputs.
$Accuracy=\frac{TP+TN}{TP+FN+TN+FP}=\frac{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|+|\{p^{\prime}\in\mathcal{P}_{b}:detect(R^{\prime}_{p})=0\}|}{|\mathcal{P}_{a}|+|\mathcal{P}_{b}|}$
where $\mathcal{P}_{a}$ and $\mathcal{P}_{b}$ are the datasets of attack inputs and benign inputs. $\{p\in\mathcal{P}_{a}:detect(R_{p})=0\}$ and $\{p^{\prime}\in\mathcal{P}_{b}:detect(R^{\prime}_{p})=1\}$ separately represents attack inputs in $\mathcal{P}_{a}$ that are filtered by the detector and benign inputs in $\mathcal{P}_{b}$ that pass through the detector.
Precision calculates the number of correctly detected and filtered attack samples among all detected and filtered samples.
$Precision=\frac{TP}{TP+FP}=\frac{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|}{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|+|\{p\in\mathcal{P}_{b}:detect(R_{p})=1\}|}$
where $\{p\in\mathcal{P}_{b}:detect(R_{p})=1\}$ represents benign inputs that are incorrectly filtered by the detectors. Low precision indicates that the detector misjudges and blocks many benign inputs.
Recall assesses the number of correctly detected and filtered attack samples among all attack samples in the dataset.
$Recall=\frac{TP}{TP+FN}=\frac{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|}{|\mathcal{P}_{a}|}$
Low recall indicates that a considerable number of attack samples have bypassed the detector.
Implementation. JailGuard generates $N=8$ variants for each input. For the baseline SmoothLLM that also needs to generate multiple variants, we have recorded the detection performance of each method in SmoothLLM when producing 4 to 8 variants and display the best detection results (i.e., the highest detection accuracy) each method achieves in Table 3. For text inputs, the probability of selecting and executing the replacement, insertion, and deletion operation on each character is $p=0.005$. Notably, for important sentences, this probability is increased to 5 times the usual, resulting in a value of 0.025. JailGuard use the string ‘[Mask]’ to replace and insert. The LLM systems and applications we used on text and image inputs are the GPT-3.5-Turbo-1106 and MiniGPT-4 respectively. It is important to note that in real-world scenarios, JailGuard should be integrated and utilized as part of the LLM system and application workflow to thwart potential attacks, which means that JailGuard performs detection from the perspective of developers. Consequently, it should have access to the underlying interface of LLMs, enabling it to query multiple variants in a batch and obtain multiple responses simultaneously. In our experiments, we simulate this process by making multiple accesses to the LLM system’s API. Our framework is implemented on Python 3.9. All experiments are conducted on a server with AMD EPYC 7513 32-core processors, 250 GB of RAM, and four NVIDIA RTX A6000 GPUs running Ubuntu 20.04 as the operating system.
6.2. RQ1: Effectiveness of Detecting Attack
Fig. 5. Comparison of Different Methods’ Results (Red marks baselines and blue marks JailGuard’s mutators and policies. The upper right indicates the best results.)
Experiment Designs and Results. To demonstrate the effectiveness of JailGuard in detecting and defending LLM attacks, we evaluate mutators and combination policies in JailGuard and all baselines on our whole text and image datasets. The results on text and image inputs are separately shown in Table 3 and Table 4. The rows of ‘Baseline’ show the detection results of four baselines on text and image inputs, and ‘JailGuard’ rows correspond to the detection results of applying different mutation strategies in JailGuard. The default combination policy is marked in italics. The row ‘Average’ shows the average result of baselines and JailGuard. The names of JailGuard’s mutators and baselines refer to § 4.2 and § 6.1. We use ‘*’ to mark the baseline method with the highest accuracy, which has the best performance in identifying both attack and benign samples. In addition, we bold the results of those mutators in JailGuard which achieves higher accuracy than that of the best baseline. In addition, Fig. 5 uses two scatter plots to compare the detection results between baselines and JailGuard on text and image modalities. The X-axis is the recall and the Y-axis is the precision. Blue dots indicate the results of mutators and policies in JailGuard and red dots mark the baselines. The methods or mutators represented by each dot are detailed at the top of the table.
Table 3. Comparison of Attack Mitigation on Text Inputs (* Marks The Highest Accuracy of Baseline. Bold Marks Results That Outperform the Best Baseline. Blue Marks the Best Results of JailGuard)
Method Acc. (%) Pre. (%) Rec. (%)
Baseline Content Detector 60.41 50.52 49.48
SmoothLLM-Insert 73.89 83.28 43.45
SmoothLLM-Swap 74.33* 83.09 44.98
SmoothLLM-Patch 72.53 80.94 40.98
In-Context Defense 73.09 80.92 42.83
Paraphrase 68.63 74.45 32.85
Perplexity-based Detection 43.23 41.29 99.43
Data Prompt Isolation 68.03 74.26 30.73
LLM-based Detection 72.55 60.88 87.78
Prompt Learning 70.25 75.84 37.60
Self-reminder 73.17 82.08 42.13
Average 68.19 71.70 50.11
JailGuard Random Replacement 80.95 75.16 78.23
Random Insertion 81.31 80.59 70.18
Random Deletion 82.40 79.57 75.35
Punctuation Insertion 81.40 84.34 65.70
Synonym Replacement 75.21 65.77 79.30
Translation 80.93 72.84 83.43
Targeted Replacement 82.02 74.27 84.23
Targeted Insertion 84.73 82.04 79.15
Policy 86.14 80.58 86.10
Average 81.68 77.24 77.96
Analysis. The results in Table 3 and Table 4 demonstrate the effectiveness of JailGuard in detecting LLM prompt-based attacks across different input modalities. JailGuard achieves an average detection accuracy of 81.68% on text inputs and 79.53% on image inputs with different mutators, surpassing the state-of-the-art baselines, which have an average accuracy of 68.19% on text inputs and 66.10% on image inputs. Remarkably, all mutators and policies implemented in JailGuard surpass the best baseline, with their results highlighted in bold. In addition, JailGuard achieves an average recall of 77.96% on text inputs and 77.93% on image inputs, which is 1.56 and 3.50 times the average result of baselines (50.11% and 22.25%), indicating its effectiveness in detecting and mitigating LLM attacks across different modal inputs. While excelling in attack detection, JailGuard also reduces FPs and separately improves the averaged precision by 5.54% and 1.19% on text and image inputs.
On the text dataset, the mutators and policy in JailGuard achieve an average accuracy and recall of 81.68% and 77.96%, which is 13.49% and 27.85% higher than the average results of the baselines. The best baseline (i.e., the ‘swap’ method of SmoothLLM) achieves the highest accuracy of 74.33% and recall of 44.98%. In comparison, different mutation strategies in JailGuard improve the accuracy of the best baseline by a factor of 1.18%-15.89% and improve its recall by 46.06%-113.65%. More specifically, the Random Insertion and Random Deletion mutators achieve the best accuracy of 82.40% and 81.31% among random mutators. The word-level and sentence-level mutators Synonym Replacement and Translation achieve the worst accuracy, namely 75.21% and 80.93%. Our analysis of their results shows that when creating variants, synonym replacement and translation can cause subtle changes in the semantics of words and sentences, leading to more false positives of benign cases. Although these two methods have good detection results on attack inputs (i.e., high recall), the increase in false positives limits their overall performance. In addition, all targeted mutators achieve much better results than their random version. Targeted Replacement and Targeted Insertion separately achieve the detection accuracy of 82.02% and 84.73%, improving the accuracy of 1.07% and 3.42% compared to Random Replacement and Random Insertion. Further analysis of their detection results reveals that the advantage of targeted mutators lies in detecting attacks with long texts and complex templates These attacks often use templates to construct specific scenarios and role-playing situations. The targeted mutators can identify the key content through word frequency and apply additional disturbances, thereby interfering with these attack samples and achieving better detection results. This observation is further confirmed by the ’Template’ column in Fig. 6. In addition, the combination policy in JailGuard further achieves the highest accuracy of 86.14% (marked in blue in Table 3), which illustrates the effectiveness of the mutator combination policy. We further study the impact of the probability in the built-in policy on detection results in § 6.4.
Table 4. Comparison of Attack Mitigation on Image Inputs (* Marks The Highest Accuracy of Baseline. Bold Marks Results That Outperform the Best Baseline. Blue Marks the Best Results of JailGuard)
Method Acc. (%) Pre. (%) Rec. (%)
Baseline Self-reminder 61.50 60.87 10.50
ECSO Detection 70.70* 82.42 34.00
Average 66.10 71.65 22.25
JailGuard Horizontal Flip 79.60 72.90 78.00
Vertical Flip 81.00 74.42 80.00
Random Rotation 80.20 74.28 77.25
Crop and Resize 77.80 72.14 72.50
Random Mask 78.80 71.66 77.75
Random Solarization 77.70 69.71 78.25
Random Grayscale 81.10 76.18 76.75
Gaussian Blur 79.50 73.49 76.25
Colorjitter 76.90 69.07 76.50
Random Posterization 79.30 73.37 75.75
Policy 82.90 74.00 88.25
Average 79.53 72.84 77.93
On the image dataset, the baseline methods achieve an average accuracy of 66.10% and recall of 22.25%. The best baseline, ECSO detection, achieves an accuracy of 70.70% and a recall of 34.00%, illustrating the limitations of baselines in detecting attacks on image inputs. In contrast, the mutation strategies in JailGuard have achieved an average accuracy of 79.53% and recall of 77.93%, which far exceeds the results of baselines. The mutators and policy improve the best detection accuracy of baselines by a factor of 8.77%-17.26%, and the improvement on recall is even more significant, that is, 113.24%-159.56%. The policy in JailGuard combines the mutators Random Rotation, Gaussian Blur and Random Posterization, further achieving the detection accuracy and recall of 82.90% and 88.25%. It improves the results of the best baseline by 12.20% and 54.25%, demonstrating the detection effectiveness of JailGuard’s policy.
In addition, Fig. 5 intuitively demonstrates the advantages of JailGuard in attack detection compared to the baselines. We can observe that JailGuard (blue) achieves significantly better results than baselines (red), and the corresponding dots are distributed in the upper right corner, indicating high precision and recall in detection.
Answer to RQ1: All mutation strategies in JailGuard can effectively detect LLM prompt-based attacks on text and image inputs, surpassing state-of-the-art methods in detection accuracy. JailGuard achieve an average accuracy of 81.68% and 79.53% on image and text datasets, respectively. For the single mutators, targeted mutators can achieve better detection results than their random versions, improving accuracy by 1.07% and 3.42%. Moreover, the combination policies in JailGuard further improve the detection accuracy to 86.14% and 82.09% on text and image inputs, significantly outperforming state-of-the-art detection methods by 11.81%-25.73% and 12.20%-21.40%, demonstrating the effectiveness of the default combination policy in JailGuard.
6.3. RQ2: Effectiveness of Detecting Different Kinds of Attacks
Experiment Designs and Results. To demonstrate the effectiveness and generalization of JailGuard in detecting various LLM attacks, we analyze the detection accuracy of the defense methods on each attack method and display the results as heat maps, as shown in Fig. 6 and Fig. 7. Each column represents an LLM attack method, which is collected in our dataset, as mentioned in § 5. Fig. 6 shows the detection results of samples in the text dataset. The first seven columns are the detection accuracy on different jailbreaking attack samples, the following five columns indicate the results on hijacking attack samples, and the last column shows the detection results on benign samples. Fig. 7 shows the detection results on the image dataset. The first three columns are the detection results of the typographic attacks on stable diffusion images, typographic attacks on blank images, and jailbreaking attacks based on adversarial perturbations. The last column of the two figures is the detection accuracy of benign samples. A bluer color on the heat maps signifies higher accuracy in detecting a specific input type, otherwise, it means that the method struggles to identify that type of input. The blank row on the heat maps separates the results of baseline methods (upper part) from JailGuard (lower part). Results for targeted mutators and the default combination policies in JailGuard are highlighted in italics and bold.
Fig. 6. Comparison of Different Methods’ Results on Text Inputs
Fig. 7. Comparison of Different Methods’ Results on Image Inputs
Analysis. The experiment results illustrate the effectiveness and generalization of JailGuard’s mutators, especially for the targeted mutators and policies, in detecting various attacks. From Fig. 6 and Fig. 7, we observe that most baseline methods struggle to detect attack samples with different attack targets. The jailbreak defense methods (e.g., SmoothLLM and In-Context Defense) usually leverage jailbreaking cases and keywords to detect attacks. As a result, they can hardly provide effective detection for hijacking attacks with unknown attack targets. Although perplexity-based detection and LLM-based detection can effectively block most attack samples, they introduce a large number of false positives, allowing only 5.77% and 62.40% of benign samples to pass, which is significantly lower than other methods. Furthermore, even for jailbreaking attacks, there is substantial variability in baseline detection effect for samples generated by different attack methods. For example, the SmoothLLM with ‘insert’ method only has a detection accuracy of 62.80% and 55.56% on the jailbreaking attack ‘Jailbroken’ and ‘Parameter’, which is much lower than the accuracy on other attacks (e.g., 90.32% on GPTFuzz). Similar observations can be made on the image dataset, where ECSO detection accuracy varies from 18.00% to 48.00% across different attacks. In contrast, the mutators and policies JailGuard can effectively identify various prompt-based attacks regardless of their attack targets, consistently achieving over 70% accuracy on benign samples. It indicates that JailGuard can overcome the existing limitations and exploit the divergence of variants to provide general, effective detection for various LLM prompt-based attacks.
Moreover, we also observe that different types of mutators exhibit significantly varied performances in detecting different attacks. Among the random mutators, character-level mutators (the first four rows in the lower part of Fig. 6) can hardly achieve high detection results in template-based jailbreaking attacks with lengthy content. The root cause of their poor detection effect lies in the nature of character-level perturbations, which are randomly applied and fail to affect the overall semantics of the template. For instance, Punctuation Insertion randomly inserts punctuations in the text, making it ineffective in interfering with jailbreaking attacks featuring long texts, as shown in Table 1. Consequently, its detection accuracy for attack inputs generated by GPTFuzz, Template, etc. is the lowest among all mutators. In contrast, word-level and sentence-level mutators have achieved high detection accuracy on these jailbreaking attacks with long texts. Synonym Replacement achieves an accuracy of 96.84% on TAP attacks, and Translation achieves an accuracy of 90.31% and 88.84% on Template and Jailbroken attacks. Unfortunately, excessive modifications to words and sentences can disrupt the semantics of benign samples, leading to false positives. Therefore, the accuracy on benign inputs of these two mutators is only 72.48% and 79.27%. As an improvement over random mutators, targeted mutators and policy in JailGuard can achieve better detection results on different attack samples, especially for long jailbreaking attacks and various injection attacks. Notably, the combination policy achieves accuracies exceeding 70.00% on ten attacks and 86.17% accuracy on benign samples, representing the best overall performance among all baseline methods and mutators. Additionally, the policy in JailGuard also achieves the best overall detection results on the image dataset, as shown in Fig. 7.
Fig. 8. A Case Study of Detecting ‘Jailbroken’ Attack
Case Study. We provide a case in Fig. 8 to understand and illustrate the root cause of the effect difference between JailGuard and the best baseline SmoothLLM with ‘swap’ method on specific attacks, such as Jailbroken attacks. The upper part shows the detection process of SmoothLLM and the lower part of Fig. 8 shows the detection of JailGuard combining the mutators Punctuation Insertion, Targeted Insertion, and Translation. We highlight the mutator Targeted Insertion in this case. Fig. 8.a) provides a real example of the Jailbroken attack in our dataset. Jailbroken attacks use specific content or tasks to construct scenarios and conduct role-playing, thereby deceiving the defense mechanism of the LLM system and achieving attack. SmoothLLM randomly swaps 10% characters to infer these attack inputs as much as possible. However, the impact of these perturbations is minimal. Among the eight perturbed inputs, only one attack fails, with its response containing refusal keywords, highlighted in red texts in the upper part of Fig. 8.c). Therefore, in the aggregation step in the upper part of Fig. 6.d), since most results do not contain refusal keywords, according to its aggregation principle, this input sample is incorrectly judged as a benign sample, resulting in a false negative.
In contrast, JailGuard effectively identifies and mitigates such attacks Firstly, the Targeted Insertion mutator effectively finds the important sentences of the input (marked in underlined italics) and purposefully inserts many masks to achieve interference, as shown in the lower part of Fig. 8.b). For LLM system responses shown in Fig. 8.c), JailGuard calculates their semantic similarity and divergence in Fig. 8.d) and then detects this attack based on the threshold $\theta$. In addition, even if the targeted mutator cannot effectively detect attack samples, the policy combines character-level and sentence-level mutators, effectively interfering with various attacks. Even in the situation that only one variant attack fails, since the semantics of the failed response are completely different from others, JailGuard can effectively detect it based on divergence, which makes it achieve high detection accuracy on complex attacks like Jailbroken. This case highlights the effectiveness of the targeted mutators and combination policy in JailGuard over baseline methods in detecting complex and scenario-based LLM attacks.
Answer to RQ2: Compared to baselines, the mutators and policies in JailGuard exhibit better generalization ability across different types of attacks and can effectively distinguish between prompt-based attacks and benign inputs. Moreover, the combination policy in JailGuard demonstrates stronger generalization than single mutators across various attacks. Specifically, the text mutator combination policy has achieved over 70.00% detection accuracy on 10 types of attacks, while maintaining a benign input detection accuracy of 86.17%.
6.4. RQ3: Ablation Study
Experiment Designs and Results. The mutator combination strategy in JailGuard achieves the best detection accuracy among all baselines and mutation strategies on text and image datasets. It leverages two modules to effectively detect prompt-based attacks in LLM systems and applications, which are the mutator combination policy and the divergence-based detection. To understand their contribution, we conduct an ablation experiment on the text inputs. The results, shown in Table 5, record the accuracy, precision, and recall of each method. Firstly, we implement three random policies and record their detection results to illustrate the effectiveness and contribution of JailGuard’s built-in policy. The first policy uses random probability and the same mutator pool as the built-in policy (Row ‘Random 1’), the second one uses the same probability as the built-in policy and a random mutator pool (i.e., Random Insertion, Synonym Replacement and Random Deletion, shown in Row ‘Random 2’ ), and the third one randomly select mutators from all text mutators (Row ‘Random 3’). Additionally, to understand the contribution of divergence-based detection, we substitute the divergence-based detection in JailGuard with two keyword detection methods and use the variants generated by the built-in policy to detect attacks. Their results are shown in the last two rows of Table 5. The first detection method randomly picks one variant response and uses refusal keyword detection to detect attacks (Row ‘Random Selection + Keywords’). The second detector leverages the aggregation method in SmoothLLM to get the final responses from the variant responses and then uses keywords to identify attacks (Row ‘Aggregation + Keywords’).
Analysis. The experimental results of the ablation study illustrate the effectiveness of both combination policy and divergence-based detection in JailGuard. Firstly, altering the built-in policy will degrade the detection effect. The results of Table 5 demonstrate that employing a random mutator pool or probabilities in the policy degrades detection performance, sometimes even falling below the accuracy achieved with a single operator. Random Policy 1, which uses unoptimized probabilities to select mutators from the pool, reduces the detection accuracy by 3.17% compared to the built-in policy in JailGuard. Random policies 2 and 3 that use random mutator pool achieve an accuracy of 79.18% and 82.31% respectively, which are 6.96% and 3.83% lower than the original policy. Especially for policy 2, random selection from the mutators with poor detection performance even causes the precision to drop by 10.57%.
Table 5. Ablation Study on JailGuard (Bold marks the original results of the policy in JailGuard)
Method Acc. (%) Pre. (%) Rec. (%)
JailGuard-Policy 86.14 80.58 86.10
- Random Policy 1 82.97 75.53 84.95
- Random Policy 2 79.18 70.01 83.88
- Random Policy 3 82.31 75.07 83.50
- Random Selection + Keywords 72.64 82.05 40.45
- Aggregation + Keywords 73.82 85.22 41.80
In addition, the divergence-based detection in JailGuard has an important contribution to attack detection, especially in eliminating FNs and improving recall. As shown in Table 5, using the keywords detection methods to replace the divergence-based detection in JailGuard leads to significant degradation in detection effects. Using aggregation and keyword detection will reduce the detection accuracy from 86.14% to 73.82%. Randomly selecting responses leads to even more severe performance degradation, with accuracy falling to 84.33% of the original accuracy. Worse still, the recall of random selection and keyword detection is only 40.45%, indicating that more than half of the attack samples could bypass the detection. Our analysis of the detection results shows that the keyword detection overlooks many attack examples, particularly hijacking prompt injection attacks, and cannot provide effective defense for various attacks. This observation is consistent with our findings in § 6.3.
Answer to RQ3: Both the built-in mutator combination policy and divergence-based detection framework of JailGuard have a significant contribution to achieving effective detection. Modifying the combination policy or the divergence-based detection leads to performance degradation, potentially allowing over 50% attack samples to evade detection.
6.5. RQ4: Impact of Variant Amount
Fig. 9. Impact of Variants‘ Number (Budget) on Detection Results
Experiment Designs and Results. JailGuard leverage the mutation strategies to generate $N$ variants and compute the divergence in the corresponding responses. To understand the impact of different values of $N$ (i.e., different LLM query budget) on detection results, we evaluate the detection effectiveness of different mutators and policies when generating 4/5/6/7/8 variants, and record accuracy and recall on the image and text dataset, as shown in solid lines with different colors Fig. 9. The bold red line marks the result of the default combination policy in JailGuard. In addition, the number of variants also affects the detection effect of the baseline SmoothLLM (Robey et al., 2023). We have run the three methods (i.e., ‘insert’, ‘swap’, and ‘patch’) of SmoothLLM on 4 to 8 variant budgets. Then we record the best results achieved by SmoothLLM in the whole experiment, that is, when using the ‘swap’ method and generating 8 variants, the accuracy reaches 74.33% and when generating 4 variants with ‘swap’, SmoothLLM has the highest recall of 47.98%, as shown in the dashed purple lines in Fig. 9.a).
Analysis. We can observe that decreasing the number of variants (i.e., LLM query budget) generally degrades the detection results of the mutators and policies. Taking the combination policy of JailGuard as an example, when the number of variants drops from 8 to 4, the detection recall on the text drops from 86.10% to 82.25%, and it is more obvious on the image, from 88.25% to 72.50%. However, such a decrease is not obvious in detection accuracy, which only drops 1.17% and 5.30% on text and image datasets respectively. Our analysis shows that benign samples are less likely to exhibit large divergence when producing fewer variants, leading to fewer false positives and mitigating the reduction in detection accuracy. Notably, for the single mutator Synonym Replacement which has the lowest detection precision among all mutators, as shown in Table 3. A reduction in the query budget actually increases its detection precision from 65.77% to 69.57%, and its detection accuracy from 75.21% to 76.94%. Moreover, the detection results of any mutator or policy under any query budget outperform the best results achieved by the baseline SmoothLLM (the dashed purple lines in Fig. 9.a)), which illustrates the potential of JailGuard for effective attack detection even under low query budget.
In actual deployment scenarios, JailGuard accesses LLM to batch process and infer the input variant, which leads to additional memory overhead. Our simulations on MiniGPT-4 show that a single set of inputs (one image and one corresponding instruction) increases the memory overhead by 0.49GB, which is equivalent to 3.15% of the LLM memory overhead (15.68GB). If the LLM query budget is set to $N=8$, the memory overhead of JailGuard to detect jailbreaking attacks is 3.95GB, which is 25.20% of the memory overhead of LLM itself. Although the runtime overhead of JailGuard is acceptable, considering that resources may be limited in LLM system application and deployment scenarios, performing effective attack detection with lower overhead has great significance. Considering that under different settings of budget $N$, JailGuard can consistently achieve detection results far exceeding the baseline under different budget settings and the detection effect improves with the increase of $N$, we recommend using $N=8$ as the default number of variants to achieve the best detection effect across different attacks and using $N\in[4,6]$ to achieve the balance between the detection effect and runtime overhead in resource-constrained scenarios.
Answer to RQ4: Reducing the query budget and the number of variants results in a slight degradation in JailGuard detection accuracy and recall. Even so, in the process of reducing the budget, the detection effect of the mutators and policies in JailGuard is always better than the best results of SmoothLLM, indicating the potential of JailGuard to detect prompt-based attacks in low-cost scenarios. Consequently, JailGuard generates 8 variants by default to obtain the best detection effect. When the LLM query budget is limited, users can choose to generate 4 to 6 variants to obtain a balance of efficiency and effectiveness.
7. Related work
LLM Attack And Defense Supplement to the prompt-based attack methods in § 2, researchers proposed other methods to automatically generate jailbreak and hijacking prompts (Niu et al., 2024; Geiping et al., 2024; Deng et al., 2023b; Liu et al., 2024b; Zhang et al., 2024; Andriushchenko et al., 2024; Deng et al., 2023a; Shen et al., 2024). Geiping et al. (Geiping et al., 2024) construct misleading, misinformation, and other non-jailbreaking attack instructions based on the existing jailbreak attacks (Zou et al., 2023). Unfortunately, we cannot find available open-source code or datasets of their attacks. Researchers also pay attention to other aspects of LLM security, e.g., backdoor attack (Abdelnabi et al., 2023; Xu et al., 2023; Huang et al., 2023b), privacy stealing attack (Sha and Zhang, 2024). We focus on the defense of multi-modal prompt-based attacks in this paper, which uses prompts as the carrier and does not require finetuning or modification of the target LLMs and systems.
To defend against LLM attacks, in addition to the baselines in § 6.1, researchers have proposed other methods (Kumar et al., 2023; Chen et al., 2024b; Pi et al., 2024; Zhou et al., 2024; Wu et al., 2024; Zhou et al., 2024). Kumar et al. (Kumar et al., 2023) designed a detection method that splices the input text and applies a safety filter on all substrings to identify toxic content. However, this method will have a significant overhead on a long input prompt. Similar to Self-reminder (Xie et al., 2023), Self-defend (Wu et al., 2024) uses system prompts to ask LLMs to self-check whether the given input is an attack input. Unfortunately, we cannot find available open-source implementation. In this paper, we compare JailGuard with 12 state-of-the-art open-sourced detection and defense methods.
Adversarial Attack and Defense in DNNs. White box attacks assume the attacker has full knowledge of the target model, including its architecture, weights, and hyperparameters. This allows the attacker to generate adversarial examples with high fidelity using gradient-based optimization techniques, such as FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), Square Attack (Andriushchenko et al., 2020). AutoAttack (Croce and Hein, 2020) has been proposed as a more comprehensive evaluation framework for adversarial attacks. Recently, researchers have also been exploring the use of naturally occurring degradations as forms of attack perturbations. These include environmental and processing effects like motion blur, vignetting, rain streaks, varying exposure levels, and watermarks (Gao et al., 2022; Guo et al., 2020; Jia et al., 2020; Tian et al., 2021; Hou et al., 2023).
Adversarial defense can be categorized into two main types: adversarial training and adversarial purification (Nie et al., 2022). Adversarial training involves incorporating adversarial samples during the training process (Goodfellow et al., 2014; Madry et al., 2017; Athalye et al., 2018; Rade and Moosavi-Dezfooli, 2021; Ding et al., 2018), and training with additional data generated by generative models (Sehwag et al., 2021). On the other hand, adversarial purification functions as a separate defense module during inference and does not require additional training time for the classifier (Guo et al., 2017; Xu et al., 2017; Sun et al., 2019; Ho and Vasconcelos, 2022).
8. Discussion
JailGuard EnhancementJailGuard requires several additional LLM queries to generate variant responses and detect attacks. Even if it can generate a smaller number of variants (i.e., $N=4$), this extra runtime overhead is still unavoidable. Moreover, fewer variants lead to a degradation in detection results. Developing a more effective mutation strategy that maintains high detection accuracy with a lower query budget is a critical area for future research. One possible solution is to utilize speculative small models (e.g., GPT-3) to generate variant responses, thus reducing queries costs. ❷ The heuristic benign instructions are the main causes of the false positives in JailGuard (e.g., ‘Suggest some names for a friendly group in telegram’). They have no clear answers and their responses are prone to high divergence, which significantly contributes to false positives in detection. Identifying such heuristic benign questions and mitigating false positives in attack detection is a crucial challenge for enhancing JailGuard. One potential approach involves designing an AI-based filter to automatically filter out these heuristic benign inputs. ❸ JailGuard currently implements 18 mutators and a set of mutator combination policies for the inputs on text and image modalities. With the development of MLLMs, audio input is becoming another important modality (e.g., GPT-4o (OpenAI, 2024)). Existing work (Wu et al., 2023; Abdullah et al., 2021) has pointed out the characteristics of poor robustness and transferability of audio adversarial attacks. Based on such observations, although there are currently no relevant MLLM audio attack methods and datasets, the detection metric in JailGuard (i.e., divergence) still has feasibility in detecting audio attacks. How to design mutators for audio attacks will be a potential future direction. Our framework and dataset will be continuously updated and any new appearing attacks will be further collected and evaluated. You can find the latest information on our website (our, 2024).
Diverse LLM Attacks Detection ❶ As an emerging research field, the security of large models has received widespread attention from researchers and industry. It is significant to add more types of attack inputs (e.g., data poisoning (Yan et al., 2023) and backdoor (Abdelnabi et al., 2023; Xu et al., 2023), and misinformation (Geiping et al., 2024) ) and build a comprehensive and universal benchmark for LLM defense. ❷ Our detection method fundamentally leverages the inherent non-robustness of attacks. Consequently, the vulnerabilities introduced by data poisoning and model backdoors, which also exhibit this non-robustness, could potentially be identified by our detection framework. A crucial future direction involves designing defense methods that are both effective and efficient, capable of generalizing across various types of attack inputs. Successfully achieving this would significantly enhance the deployment and application of trustworthy Language Model (LM) systems, contributing to their overall reliability and security.
9. Conclusion
In this paper, we propose JailGuard, a universal detection framework that detects both jailbreaking and hijacking attacks for LLM systems on both image and text modalities. To comprehensively evaluate the detection effect of JailGuard, we construct the first comprehensive LLM prompt-based attack dataset, covering 15 jailbreaking and hijacking attacks and 11,000 items of data on image and text modalities. Our experiment results show that JailGuard achieves the best detection accuracy of 86.14%/ 82.90% on text/image inputs, significantly outperforming state-of-the-art defense methods by 11.81%-25.73% and 12.20%-21.40%.